# VectorLiteRAG: Latency-Aware and Fine-Grained Resource Partitioning for Efficient RAG

### v2

@@ VIII-A Performance Model and Hit Rate Estimator
Figure 10 evaluates the accuracy of VectorLiteRag’s performance model. The right panel compares the predicted and actual minimum hit rates within each batch. The close alignment confirms that our Beta-distribution-based approximation reliably captures caching effectiveness.
As expected from order statistics, the minimum hit rate declines rapidly with increasing batch size, with the rate of decline flattening in the larger batch region.
@@
The left panel compares the predicted latency of hybrid index search against the measured latency. While the predictions follow the same trend, there is a consistent offset between the two. This discrepancy primarily stems from the dispatcher’s early query handling, which we discuss further in Section VIII-E. Despite this, the model serves as a reliable upper bound for search latency and is sufficiently accurate for guiding the partitioning algorithm.
@@
Figure 10: Comparison of measured versus estimated values from VectorLiteRAG’s performance model. Left: Predicted and actual search latency across batch sizes. Right: Estimated and measured minimum (tail) hit rates within a batch. The model accurately tracks trends and serves as a reliable upper bound for latency-aware partitioning.
Figure 11: End-to-end latency of the full RAG pipeline under varying arrival rates for different datasets. Compared systems include CPU-only retrieval (blue), dedicated GPU for retrieval (green triangle), all-GPU (green square), and vLiteRAG (orange) just as SLO attainment. vLiteRAG consistently maintains lower or comparable latency under high load conditions.
VIII-B SLO Attainment
@@
Achieving service level objectives for TTFT across a wide range of input traffic is the primary goal of VectorLiteRAG. Figure 9 presents the SLO attainment curves across all nine combinations of vector databases and LLMs evaluated. In each subplot, the horizontal dashed line marks the 90th percentile latency target, while the vertical dashed line indicates the standalone LLM throughput. All experiments use an on-demand dynamic batching framework, where retrieval requests are served as soon as the previous search completes. This batching mechanism scales throughput with arrival rate by adapting batch sizes dynamically.
@@
Across all configurations, VectorLiteRAG consistently sustains the extended SLO budget ($SLO_{\text{LLM}}+SLO_{\text{Search}}$, defined in Table I) over the widest input rate ranges among evaluated baselines. CPU-based fast scan can support relatively high request per second (RPS) rates, its limited per-request performance leads to consistent SLO violations even under light traffic. As arrival rate increases, batch sizes grow (up to 9–10 under >>40 RPS), incurring high search latency and poor tail response.
@@
For small vector databases and under light loads, all-GPU shared configurations can satisfy SLOs over wide traffic ranges. However, as arrival rate approaches the reduced LLM throughput, due to index memory contention latency increases sharply. Although VectorLiteRAG is subject to this limitation as well, its optimized partitioning algorithm extends the SLO-attainable region nearly up to the standalone LLM throughput limit.
@@
Figure 12: TTFT latency breakdown for systems using the Wiki-All and ORCAS 1K indexes with Qwen3-32B. Each group shows results from four configurations: DED-GPU, All-GPU, VectorLiteRAG, and CPU-Only (from left to right). Bars are stacked to show the contribution of queuing delay, vector search latency (colored segments), and LLM prefill latency (grey)
VIII-C End-to-End Latency
@@
In the low-traffic regime, GPU resource contention is minimal, except in cases where dedicated GPUs are allocated solely for retrieval. However, under high traffic and with large vector databases, contention becomes significant. This is evident in the more than 2× increase in end-to-end latency observed in the all-GPU baselines for ORCAS 2K with Llama3-8B and Qwen3-32B. Although Llama3-70B involves more intensive computation, its low throughput ceiling causes TTFT to diverge before retrieval-induced interference becomes the dominant factor.
In contrast, VectorLiteRAG matches CPU-based retrieval in end-to-end latency. This demonstrates that its partitioning strategy and distributed pipeline effectively minimizes interference by carefully limiting GPU memory and compute usage for retrieval, thereby preserving LLM generation performance.
Figure 13: Comparison with HeterRAG. While HeterRAG achieves lower TTFT initially, its latency increases beyond the throughput-bound caching limit. VectorLiteRAG  is configured with $SLO_{search}=400$ms.
VIII-D Comparison with HeterRAG
In this section, we compare VectorLiteRAGwith HeterRAG [10], that also exploits skewed cluster access patterns in RAG pipelines. While both systems adopt tiered caching strategies for vector index, their partitioning principles and system targets differ fundamentally.
HeterRAG selects GPU-resident clusters by identifying the maximum KV cache size that can sustain the throughput of the slower stage, either the LLM or the retriever. While this approach is simple and throughput-aware, it is inflexible to latency constraints that are critical for real-time serving. In configurations where the LLM stage exhibits lower peak throughput than retrieval, as in our main evaluations, HeterRAG allocates the entire GPU memory to LLMs, leaving vector search to be performed on the CPU. As stated in their paper, the benefit of HeterRAG’s scheme primarily applies to scenarios where retrieval becomes extremely heavy.
To provide a fair comparison, we replicated the HeterRAG setting by building an IVF index with the number of clusters set to $\sqrt{N_{\text{vector}}}$ and measured retrieval throughput using batch sizes below 64. At $\text{{nprobe}}=256$, CPU-only retrieval achieves 35 RPS at 0.94 NDCG@50. nprobe in our system is increased accrodingly to 6144 to match this accuracy level. We implemented both systems using their respective partitioning policies. Since HeterRAG does not support distributed retrieval, we applied their GPU caching scheme using IndexIVFShard without our optimized pipeline.
Figure 13 summarizes the results. HeterRAG places 73% of the index clusters in GPU memory, whereas VectorLiteRAG identifies a partitioning point at 31.5% of clusters under a 400ms SLO constraint. HeterRAG achieves lower retrieval latency under low traffic, but its operable region is limited as input rates increase. VectorLiteRAG maintains latency around the specified constraint across a wider range of traffic levels. In addition, thanks to efficient resource usage in our distributed pipeline it also achieves lower end-to-end latency overall through its optimized, distributed retrieval pipeline.
The key difference lies in how partitioning decisions are made. VectorLiteRAG enables the system operator to specify a target SLO and computes the largest feasible GPU-resident index region that meets this constraint. In contrast, HeterRAG determines partitioning solely based on balancing throughput between stages, without considering latency objectives. As a result, it lacks fine-grained latency control and may misallocate GPU resources.
VIII-E Dynamic Dispatcher
Figure 14 illustrates the effectiveness of the dynamic dispatcher in the VectorLiteRAG pipeline. By polling the scanning loop and dispatching queries as soon as they complete, the dispatcher reduces search latency by up to 16%. This optimization benefits not only the average latency, but also the tail latency, as shown in the right pane of the figure. The improvement comes from overlapping the merging and re-ranking of early-completed queries with the ongoing scanning of slower queries, thereby alleviating the overhead of merging all results in bulk at the end.
Figure 14: Left: Vector search average and Right: P90 tail latency on ORCAS 2K index with dispatcher enabled and disabled.
IX Related Works
RAG applications with iterative retrieval or multi-stage generation often exhibit semantic similarity across successive queries. Motivated by this observation, several optimization techniques have been proposed, including prefetching [22], speculative retrieval [41], and pipelined execution [14]. In contrast, our work builds upon application-agnostic, generic retrieval–generation pipelines without relying on semantic priors or intermediate signals.
RagCache [15] improves the generation throughput by managing the reuse of KV cache between tenants, focusing on scheduling on the LLM side. Hermes [33] achieves datacenter-scale throughput by offloading vector search to additional CPU nodes, at the cost of requiring disaggregated infrastructure.
Efforts like  [20, 13, 11, 30, 24] propose specialized hardware or memory centric architectures to accelerate RAG pipelines. While these approaches offer significant performance gains, they often rely on custom infrastructure, which may limit deployability in general-purpose environments.
Among prior works, HeterRAG [10] also co-locates retrieval and generation on GPU. Our work builds on this direction with an analytical model for latency and hitrate, enabling principled GPU memory partitioning under explicit SLOs. To our knowledge, VectorLiteRAG is the first solution offering fine-grained resource control for co-located RAG pipelines.
X conclusion

### v3

@@ VI-A Performance Model and Hit Rate Estimator
Figure 10 evaluates the accuracy of VectorLiteRag’s performance model. The right panel compares the predicted and actual minimum hit rates within each batch. As expected from order statistics, the minimum hit rate declines rapidly as batch size increases, and the rate of decline gradually flattens in the large-batch regime. Close alignment of two curves confirms that our Beta-distribution-based approximation reliably captures caching effectiveness.
@@
The left panel compares the predicted latency of the hybrid index search with the measured latency. While the predictions generally follow the same trend, a noticeable offset exists between the two. This deviation mainly results from the dispatcher’s early-query handling, as discussed in Section VI-E1.
Precisely capturing the dispatcher’s impact would require evaluating full order statistics to model per-request completion times, which greatly increases complexity while providing only marginal benefit. Despite these approximations, the resulting configurations perform robustly in practice, as shown in the following sections.
@@
Figure 10: Comparison of measured (solid line) vs estimated (dotted line) values from VectorLiteRAG’s performance model. Left: Search latency across batch sizes. Right: Tail hit rates within a batch.
Figure 11: Left: TTFT SLO attainment and Right: end-to-end latency of RAG pipeline under increasing arrival rates across different LLMs (columns) and datasets (rows). Our work (vLiteRAG) achieves higher SLO attainment across all regimes compared to baselines.
VI-B SLO Attainment
@@
Figure 11 presents SLO attainment curves across all nine combinations of vector databases and LLMs. In each subplot, the horizontal dashed line marks the 90th percentile latency target, and the vertical dashed line indicates the standalone LLM throughput. All experiments use on-demand dynamic batching, where retrieval requests are served immediately after the previous search completes, allowing throughput to scale with arrival rate through adaptive batch sizing.
@@
Across all configurations, VectorLiteRAG sustains the extended SLO budget ($SLO_{\text{LLM}}+SLO_{\text{Search}}$, defined in Table I) over the widest input rate ranges among evaluated baselines. CPU-based fast scan can support relatively high request per second (RPS) rates, its limited per-request performance leads to consistent SLO violations even under light traffic. As arrival rate increases, batch sizes grow (up to 9–10 under >>40 RPS), incurring high latency and poor tail response.
@@
For small vector databases and under light loads, ALL-GPU configurations can satisfy SLOs over wide traffic ranges. However, as the arrival rate approaches its reduced throughput, latency increases sharply. Although VectorLiteRAG is subject to this limitation as well, its optimized partitioning algorithm extends the SLO-attainable region nearly up to the standalone LLM throughput limit.
@@
Figure 12: TTFT breakdown for Wiki-All and ORCAS 1K indexes with Qwen3-32B. Each group shows results from four configurations. Bars are stacked to show the contribution of queuing delay, vector search latency (colored segments), and LLM prefill latency (grey)
VI-C End-to-End Latency
@@
In contrast, VectorLiteRAG matches CPU-based retrieval in end-to-end latency. This demonstrates that its partitioning strategy and distributed execution pipeline effectively minimizes interference by carefully limiting GPU memory and usage of GPU threads for retrieval, thereby preserving LLM generation performance, while maintaining latency lower than SLO requirements.
Figure 13: Comparison with HedraRAG. HedraRAG exhibits lower TTFT at low request rates, but latency increases sharply once the system exceeds its throughput limit. VectorLiteRAG  is configured with $SLO_{search}=400$ms.
VI-D Comparison with HedraRAG
We compare VectorLiteRAG with HedraRAG[11], which also exploits skewed cluster access patterns in RAG pipelines. While both systems adopt tiered caching strategies for vector indices, their partitioning principles and target objectives differ fundamentally.
HedraRAG selects GPU-resident clusters by identifying the maximum KV cache size that can sustain the throughput of the slower stage, either the LLM or the retriever. Although this approach is simple and throughput-aware, it does not account for latency constraints that are critical for real-time serving. In configurations where the LLM stage exhibits lower peak throughput than retrieval, as in 11, HedraRAG allocates the entire GPU memory to LLMs and performs vector search on the CPU. As noted in their paper, HedraRAG is most effective when retrieval becomes extremely heavy.
To enable a fair comparison, we replicate the HedraRAG setting by building an IVF index with $\sqrt{N_{\text{vector}}}$ clusters and measuring retrieval throughput using batch sizes below 64. At $\text{{nprobe}}=256$, CPU-only retrieval achieves 35 RPS at 0.94 NDCG@50; we increase nprobe to 6144 in our system to match this accuracy. Since HedraRAG does not support distributed retrieval, we apply their GPU caching scheme using IndexIVFShard without our optimized pipeline.
Figure 13 summarizes the results. HedraRAG places 73% of index clusters in GPU memory, whereas VectorLiteRAG identifies a partitioning point of 31.5% under a 400,ms SLO. While HedraRAG achieves lower retrieval latency under low traffic, its operable range narrows as input rates increase. In contrast, VectorLiteRAG maintains latency near the target constraint across a wider traffic range and achieves lower overall end-to-end latency through its distributed pipeline.
The key distinction lies in how partitioning decisions are made. VectorLiteRAG allows operators to specify a target SLO and computes the largest GPU-resident index region that satisfies this constraint, whereas HedraRAG balances throughput between stages without explicit latency objectives, which can lead to suboptimal GPU allocation.
VI-E Ablation Studies
VI-E1 Dynamic Dispatcher
Figure 14: Left: Average search latency and batch sizes. Right: P90 tail latency on ORCAS 2K index with dispatcher enabled and disabled.
Figure 14 illustrates the effectiveness of the dynamic dispatcher in the distributed VectorLiteRAG pipeline. By polling the scanning loop and dispatching queries immediately upon completion, the dispatcher reduces search latency by up to 16%, improving both average and tail latency. This gain is achieved by overlapping the merging and re-ranking of early-completed queries with the ongoing scanning of slower queries, avoiding bulk merging at the end.
Figure 14 also reports average batch sizes under varying arrival rates. With adaptive batching, requests are grouped dynamically based on current pipeline load. Since vector search has higher throughput capacity than the LLM, it absorbs higher arrival rates by increasing batch size while maintaining stable service time. In contrast, fixed or capped batch sizes lead to request backlogs and performance degradation.
VI-E2 Impact of LLM Input and Output Lengths
Figure 15 illustrates latency sensitivity to varying input and output lengths for Llama3-8B and Llama3-70B. The red dashed line denotes the combined SLO target of vector search and LLM stages, corresponding to the 1024/256 setting in Table I. For consistency, $\text{SLO}_{LLM}$ is fixed across configurations.
Longer inputs increase prefill cost, raising TTFT and shifting SLO violations to lower arrival rates as compute resources saturate. Similarly, longer outputs reduce the SLO-compliant range due to extended generation time and higher KV cache usage. Across both dimensions, VectorLiteRAG maintains serviceability over a wider range than the baselines, highlighting the robustness of its partitioning scheme.
Figure 15: Left: P90 TTFT across different input and Right: output lengths. Darker curves represent longer input/output sequences, while brighter curves correspond to shorter ones. Experiments were conducted using the ORCAS-2K index.
VI-E3 Sensitivity study on $\text{SLO}_{search}$
To evaluate the robustness of our system under varying service constraints, we test VectorLiteRAG across multiple $\text{SLO}_{search}$ targets. All plots in Figure16 use P95 TTFT as the primary metric, with P90 results additionally shown as dashed lines for VectorLiteRAG. Changing the quantile slightly expands or shrinks the SLO-compliant range; in our evaluation, the difference between P90 and P95 was at most 1 RPS.
TABLE II: SLO targets and corresponding index shard sizes.
SLO (ms) Index (GB) Param (GB) KV Cache (GB)
100 3.80 30.59 33.24
150 2.95 34.09
200 2.47 34.57
250 2.21 34.83
Table II summarizes the target SLOs and their associated memory allocations. Under relaxed SLO constraints, the latency-bounded partitioning algorithm assigns a smaller fraction of the index to GPU shards, yielding latency behavior closer to the CPU-only baseline. As the SLO becomes stricter, the latency curve moves toward the all-GPU configuration. While tighter SLOs reduce available KV-cache space and modestly shrink the operable region, VectorLiteRAG still delivers a wider SLO-compliant throughput range than the baselines, highlighting the adaptability of its partitioning strategy and the effectiveness of its execution pipeline.
Figure 16: P95 tail latency (and P90 for VectorLiteRAG) under different search-stage SLO constraints. Results are obtained using the Qwen3-32B model and the ORCAS 1K index.
VI-E4 Robustness to Hardware Capacity
Finally, we evaluate how VectorLiteRAG adapts to different hardware capacities of the system. Following the provisioning policy commonly adopted by cloud providers, which allocates additional CPU cores as more GPUs are added, we test three configurations: 4 GPUs + 32 cores, 6 GPUs + 48 cores, and 8 GPUs + 64 cores. For each configuration, we re-profile the CPU-only search latency and apply the same latency-bounded partitioning algorithm. Aside from the number of compute devices, all experiments use identical model and index setups.
The results in Figure 17 show that VectorLiteRAG sustains the target SLO across all configurations while extending the SLO-compliant throughput roughly in proportion to the number of GPUs. While the reduced memory capacity in the GPU baseline causes decoding latency to grow rapidly with scale, VectorLiteRAG effectively contains this growth, keeping decoding latency comparable to CPU-only search cases. This demonstrates that VectorLiteRAG can be readily deployed across clusters of different sizes with minimal setup effort while maintaining consistent latency behavior.
VII Related Works
RAG applications with iterative retrieval or multi-stage generation often exhibit semantic similarity across successive queries. Motivated by this observation, several optimization techniques have been proposed, including prefetching [23], speculative retrieval [44], and pipelined execution [15]. In contrast, our work builds upon application-agnostic, generic retrieval–generation pipelines without relying on semantic priors or intermediate signals. RagCache [16] improves throughput by managing KV cache reuse between tenants, focusing on scheduling and reuse optimizations on the LLM side. Hermes [35], on the other hand, scales via disaggregation by adding CPU nodes to offload vector search.
Figure 17: Left: SLO attainment (the vertical dashed line denotes bare LLM capacity) and Right: end-to-end latency measured on 4-, 6-, and 8-GPU systems. Evaluated using the Qwen3-32B model and the ORCAS 2K index.
Efforts such as [21, 14, 12, 32, 26] propose specialized hardware or memory-centric architectures to accelerate RAG pipelines. While these approaches offer significant performance gains, they often rely on custom infrastructure, which may limit deployability in general-purpose environments. Among prior works, HedraRAG [11] also co-locates retrieval and generation on GPUs. Our work builds on this direction with an analytical model for latency and hit rate, enabling principled GPU memory partitioning under explicit SLOs. To our knowledge, VectorLiteRAG is the first solution to provide fine-grained resource control for co-located RAG pipelines.
Future work may extend our approach to prefill–decode disaggregation frameworks [45, 31], where bandwidth-bound retrieval may run alongside compute-intensive prefill. This would require jointly modeling vector search and the throughput of both stages, but our framework offers a natural basis for such integration.
VIII conclusion
This paper presents VectorLiteRAG, a latency-aware orchestration framework for Retrieval-Augmented Generation (RAG) systems that explicitly manages the tight coupling between vector retrieval and LLM inference. We show that under skewed access patterns, variability in retrieval latency interacts with inference batching, causing tail effects amplification that cannot be mitigated by optimizing either stage in isolation.
VectorLiteRAG is driven by the insight that meeting strict RAG SLOs requires balancing batching behavior rather than maximizing instantaneous GPU utilization. By coordinating retrieval progress with inference scheduling, VectorLiteRAG suppresses tail cascades and sustains predictable end-to-end latency under bursty workloads. This enables SLO compliance across a substantially wider operating regime, supporting up to 1.5× higher request rates than baseline RAG systems. Across extensive evaluation, we demonstrate that these benefits generalize across latency targets, hardware configurations, and LLM input/output lengths. VectorLiteRAG further exposes explicit control knobs that allow RAG operators to trade throughput for tail latency under constrained GPU memory budgets, making it practical for real-world deployment.
IX Acknowledgment
This research was supported in part through cyber-infrastructure research resources and services provided by the Partnership for an Advanced Computing Environment (PACE) at the Georgia Institute of Technology, Atlanta, Georgia, USA. This work was partially supported by gifts from Google and AMD. The views and conclusions contained herein are those of the authors and should not be interpreted as representing the official policies or endorsements, either expressed or implied, of Georgia Tech.
-A Abstract
The artifact includes the complete source code of the core VectorLiteRAG system, together with our modified FAISS library used for hybrid CPU–GPU vector search. To ensure reproducibility of both preprocessing and evaluation, we also provide a collection of shell scripts and Python utilities that automate the full experimental workflow, including dataset preparation, index construction, performance profiling, and end-to-end RAG pipeline evaluation. These scripts are designed to reproduce all major results reported in the paper with minimal manual intervention.
All code and supporting materials are publicly available on GitHub https://github.com/sitar-lab/VectorLiteRAG-AE and Zenodo https://zenodo.org/records/18195323
-B Artifact Check-list
  • Program: Modified FAISS library, vLLM
  • Compilation: gcc-11.3, nvcc-12.1, cmake.
  • Models: Llama-3 8B, Llama-3 70B, and Qwen-3 32B.
  • Datasets: MS ORCAS and NVIDIA Wiki-All.
  • Run-time Environment: RHEL 9 with Anaconda3.
  • Hardware: Single node equipped with 8 NVIDIA L40S GPUs and 8 NVIDIA H100 GPUs.
  • Metrics: SLO attainment, end-to-end latency, vector search hit rate estimation.
  • Output: CSV logs and visualization plots.
  • Disk Space Required: $\sim$256 GB for evaluation; $\sim$1.5 TB for preprocessing and index construction.
  • Workflow Preparation Time: 40–50 hours.
  • Experiment Completion Time: 10 hours.
  • Publicly Available?: Yes.
  • Code Licenses?: CC BY 4.0
-C Description
-C1 How to access
All source cod and scripts are accessible via github repository.
-C2 Hardware dependencies
All experiments were conducted on a single L40S node or a single H100 node, each equipped with 8 GPUs. Because larger language models rely on tensor model parallelism, the H100 system is expected to provide NVLink connectivity to ensure reproducible performance. The L40S node was configured with a 32-core Intel CPU, and the H100 node with a 64-core Intel CPU. CPU core count is an important factor, as a substantial portion of the workload executes on the host processor.
-C3 Software dependencies
The evaluation environment was run on RHEL 9 (or a compatible Linux distribution) using Anaconda3. Successful compilation of the FAISS library depends on specific toolchain versions, including Python 3.10, GCC 11.3, and NVCC 12.1. Intel MKL is also required to support vectorized CPU operations.
-C4 Datasets
The Wikiall benchmark is directly downloadable. The ORCAS 1K and ORCAS 2K benchmarks require both the MS ORCAS dataset and the English Wikipedia dump, which are publicly accessible but require long preprocessing.
-D Installation and Testing
-D1 Installation
  # Create conda environment
  cd VectorLiteRAG
  conda create -n vlite -f ./scripts/env.yml
  conda activate vlite
  # Build faiss library
  git submodule update --init --recursive
  ./scripts/build.sh
-D2 Preprocessing
  # Download a samll dataset for testing
  ./database/download.sh test
  # Chunk documents and run embedding model
  ./database/encode.sh test
  # Train index and construct base IVF
  ./scripts/train.sh test
-D3 Testing
  # Run a round of short cpu search based RAG
  # Output csv files will appear under results/test
  ./scripts/test.sh
-E Experiment Workflow
We provide scripts for evaluation along with corresponding plotting utilities. The workflow is straightforward. After completing the preprocessing steps, execute the following commands from the project’s root directory:
  • (1)
    Download the datasets: ./database/download.sh
  • (2)
    Perform preprocessing and train the base indexes:
    ./database/encode.sh <dataset>
    ./scripts/train.sh <dataset>
  • (3)
    Run all experiments sequentially:
    ./scripts/runall.sh
    For running experiments individually, please refer to the repository README.
  • (4)
    After all experiments are completed, generate the figures: ./scripts/plotall.sh
    Individual plotting options are also documented in the README.
-F Evaluation and Expected Results
This artifact reproduces the primary experimental results presented in Figures 10–17 of the paper. During evaluation, latency logs are saved under results/<datasets>, and all vector indexes and their associated metadata are stored in database/<dataset>. After the evaluation completes, the provided plotting scripts generate the corresponding figures and place them in figures directory.
The reproduced results are expected to closely align with those reported in the paper, though minor variations may occur due to system-level factors. Individual evaluation runs for each data point are also supported pyand documented in the repository, enabling users to verify and assess any deviation.